OpenStudy (rebeccapink6398):

pls help!!! Will medal+fan!!! (question is in chat) @Legends @BlossomCake @~*Ashtyn*~ kelp?

OpenStudy (rebeccapink6398):

Benito solved this problem. The Martinez family is visiting relatives who live 536 miles away. The Martinez family car uses about 1 gallon of gasoline every 25 miles. They wanted to know how many gallons they would use on the way to their relatives' house. Benito said they would use about 21.4 gallons. Is Benito's answer reasonable? A. No, because 550 / 25 = 22 and since the distance is 536 miles, the answer will be slightly higher than this estimate. B. No, because 500 x 25 = 12,500, and since the distance is 536 miles, the answer will be slightly higher than this estimate. C. No, because 500 x 25 = 12,500, and since the distance is 536 miles, the answer will be slightly lower than this estimate. D. Yes, because 500 / 25 = 20 and since the distance is 536 miles, the answer will be slightly higher than this estimate.

OpenStudy (rebeccapink6398):

@Legends @BlossomCake @~*Ashtyn*~

OpenStudy (anonymous):

i got this

OpenStudy (rebeccapink6398):

XD ok

OpenStudy (anonymous):

Benito solved this problem. The Martinez family is visiting relatives who live 536 miles away. The Martinez family car uses about 1 gallon of gasoline every 25 miles. They wanted to know how many gallons they would use on the way to their relatives' house. Benito said they would use about 21.4 gallons. Is Benito's answer reasonable? A. No, because 550 ÷ 25 = 22 and since the distance is 536 miles, the answer will be slightly higher than this estimate. B. No, because 500 • 25 = 12,500, and since the distance is 536 miles, the answer will be slightly lower than this

OpenStudy (anonymous):

one of those a or b

OpenStudy (anonymous):

just make 536/25 = 21.44

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Sorry I couldn't help I am TERRIBLE at math :p lol